# What makes an image realistic?

### v3

@@ 1 Introduction
Potential applications of such functions are plentiful and include anomaly detection Ruff et al. (2021), deepfake detection (Sha et al., 2023; Pondoc et al., 2023), evaluation of generative models (Theis et al., 2016; Heusel et al., 2017), model distillation (van den Oord et al., 2018; Yin et al., 2023), neural compression (Ballé et al., 2021; Yang et al., 2023), computational photography (Fang et al., 2020), 3D synthesis (Herzog et al., 2012; Poole et al., 2023), and other content generation tasks. Unfortunately, their implementation is extremely challenging. Our ability to generate realistic data is rapidly improving (e.g., Dhariwal and Nichol, 2021) yet no reliable candidates or recipes for constructing $U$ exist in machine learning today. This is not for a lack of trying. While some progress has been made in the detection of unrealistic examples, the design of functions that are robust to optimization (for tasks involving generation) has been less successful. The latter problem is significantly harder because our function now not only has to detect a limited set of artefacts but has to anticipate any unrealistic examples an optimization might run into. Weaknesses in a function’s design often only make themselves known once we start optimizing Ding et al. (2021). Complicating the matter is the fact that the optimization depends on $U$ itself.
To give a more concrete example of the kind of tasks we are interested in, consider the following loss which naturally comes up in lossy compression. If ${\mathbf{x}}=g({\mathbf{z}})$ is the output of a neural network, we may want to find ${\mathbf{z}}$ such that
@@
In this paper we will take the view that ${\mathbf{x}}$ is realistic if it appears to have come about in a particular way, which is another way of saying that ${\mathbf{x}}$ is a plausible sample of a certain distribution $P$. What is considered realistic therefore depends on $P$. If $P$ is a distribution over natural images then most photos would qualify as realistic. While an MNIST image (LeCun and Cortes, 2010) would not be considered a realistic example of a natural image we would still consider it to be realistic if $P$ is the distibution of MNIST digits.
@@
What has been referred to as realism (e.g., Fan et al., 2018; Theis and Wagner, 2021; Careil et al., 2023) is also often referred to as perceptual quality (e.g., Blau and Michaeli, 2018; Fang et al., 2020; Salehkalaibar et al., 2023). It is therefore natural to wonder to what extent human perception should factor into its formalization. Our approach to defining realism is normative, that is, we consider how an idealized observer would judge realism. Similar to how Bayesian inference does not take inspiration from neuroscience but Bayesian decisions resemble human decisions (e.g., Knill and Pouget, 2004), we too can hope that human perception agrees with our notion of realism because humans face a similar task when deciding whether something is realistic. In Section 4.4, we will further make the case that batched universal critics not only generalize no-reference metrics and divergences—which represent the prevalent ways of formalizing realism—but are also a better model of a human observer.
@@ 2 Probability and typicality
In this section we review the two most common attempts to capture realism found in machine learning, namely those based on probability and typicality, and their failures. Similar failures of probability and typicality have been documented in the anomaly detection literature (e.g., Choi et al., 2019; Le Lan and Dinh, 2021; Osada et al., 2023) but are worth repeating as they continue to be a source of confusion.
@@ 2.1 Probability
Example 1. Consider a computer program simulating a sequence of independent and nearly unbiased coin tosses, ${\mathbf{x}}^{N}=(x_{1},\dots,x_{N})$ with $P(x_{n}=1)=0.5+\varepsilon$ for some very small $\varepsilon>0$. For reasonably large $N$, we would expect the program to output a number of 1s which is close to N/2 and we would suspect a bug if the program outputs a sequence of only 1s, yet this is the most probable sequence.
Example 1 shows that maximizing $P({\mathbf{x}})$ can lead to unrealistic examples. It also shows that $P({\mathbf{x}})$ would not detect a bug which causes a program to only output 1s. If instead we count the number of 1s, $k=\sum_{n=1}^{N}x_{n}$, and measure the probability of $k$, this bug could be detected. Does this mean we only need to find the right set of features? By ignoring some aspects of the data, we risk not detecting unrealistic examples. We might therefore conclude that we simply need to test sufficiently many features. Unfortunately this approach also runs into trouble. Consider testing whether ${\mathbf{x}}$ has 1s in even places and 0s in odd places, ${\mathbf{x}}=\texttt{0101..01}$. The probability of this sequence is approximately $2^{-N}$ so that we would reject it with high confidence if we happen to observe it. However, since all sequences have roughly the same probability, we would reject every sequence as unrealistic if we tested all features of this type.
Using densities instead of probabilities introduces an additional challenge, namely that our answer now depends on the parametrization of the data. If $P$ is an exponential distribution with rate 1, say, then values of ${\mathbf{x}}$ close to zero seem preferrable over larger values. But if we consider ${\mathbf{y}}=e^{-{\mathbf{x}}}$ instead, then all values of ${\mathbf{y}}$ would now be considered equally preferrable.
@@ 2.2 Weak typicality
and elements from this set are considered weakly typical. While other notions of typicality exist, weak typicality is the one most commonly encountered in the machine learning literature (Nalisnick et al., 2019b; Choi et al., 2019; Dieleman, 2020). The AEP implies that as $N$ increases, the probability that a randomly drawn sequence is contained in the typical set $A^{N}_{\delta}$ approaches 1 for any $\delta>0$. That is, a realistic sequence is likely to be typical. While we have stated the AEP for sequences of independent and discrete random variables, generalizations to dependent and continuous sources exist and are well known (e.g., Algoet and Cover, 1988).
The above suggests that instead of expecting the probability to be large, we should expect the negative log-probability of a realistic ${\mathbf{x}}$ to be close to the entropy—or the probability to be roughly $2^{-H[{\mathbf{x}}]}$, especially if ${\mathbf{x}}$ is high-dimensional. It therefore appears that $|-\log P({\mathbf{x}})-H[{\mathbf{x}}]|$ would be a good candidate for a measure of realism (Choi et al., 2019; Nalisnick et al., 2019b). Unfortunately, also this definition fails to quantify realism as the following examples demonstrate.
Example 2. Consider again a sequence of independent coin tosses. If the coin is unbiased, then the log-probability of any sequence is exactly the entropy, $-\log_{2}P({\mathbf{x}}^{N})=N$. In other words, in this case the probability of ${\mathbf{x}}^{N}$ under $P$ is completely uninformative and the typical set contains every sequence. Does this mean that every sequence of coin flips is realistic? Clearly, there is a sense in which the sequence 0000000000 is less realistic than 1100010100 which is not captured by weak typicality.
Example 3. As another example, consider an isotropic Gaussian distribution with density $p({\mathbf{x}})\propto\exp(-\|{\mathbf{x}}\|^{2})$. With high probability, the negative log-density of a random sample will be close to the differential entropy, which amounts to the norm $\|{\mathbf{x}}\|$ taking on a certain value. While we would expect realistic examples from our distribution to look like uncorrelated noise, optimizing for typicality will only constrain the norm. If ${\mathbf{x}}$ represents an image, our optimization will merely adjust its contrast but will not decorrelate pixels as one might hope.
@@
Further discussion of typicality can be found in Appendix A.
@@ 3 Divergences
More successful notions of realism are based on divergences between a ground-truth data distribution $P$ and a distribution $Q$ which we want to evaluate. In line with our intuitive notion of realism, if a divergence is zero, then instances of $Q$ are indistinguishable from instances of $P$, that is, we have perfect realism.
In coding theory, formalizing realism in terms of divergences (Matsumoto, 2018; Blau and Michaeli, 2019) has resulted in an improved understanding of the lossy compression problem and novel methods to solve them (e.g., Theis and Agustsson, 2021). In practical applications, generative adversarial networks (GANs; Goodfellow et al., 2014) trained with adversarial losses (which approximate divergences) significantly advanced the state of the art in the perceptual quality of generated images (e.g., Ledig et al., 2017). For the evaluation of generated images, the Fréchet inception distance (Heusel et al., 2017) has established itself as the method of choice and is based on a divergence between distributions over feature activations.
@@ 3.1 Adversarial losses
Note that this optimal critic depends on the distribution $q$ that we are trying to evaluate. In contrast, in our setting we may only have access to a single instance or a few instances drawn from $q$. Furthermore, the dependence of the critic on $q$ is responsible for optimization instabilities that are known to plague adversarial training and which we would like to avoid. In Section 4 we will discuss universal critics which are universal in the sense that they do not depend on $q$ and therefore do not require adversarial training.
@@ 3.2 Maximum mean discrepancy
Maximum mean discrepancy (MMD; Gretton et al., 2012) refers to a class of divergences which have been used for hypothesis testing as well as for generating realistic images (Li et al., 2015; Dziugaite et al., 2015). Given two sets of i.i.d. examples—${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{M}$ and ${\mathbf{x}}_{1}^{\prime},\dots,{\mathbf{x}}_{N}^{\prime}$—estimates of MMD can be used to decide whether the two sets were drawn from the same distribution. Formally, we compute
$\text{MMD}^{2}({\mathbf{x}},{\mathbf{x}}^{\prime})=\textstyle\left\|\frac{1}{M}\sum_{m}\Phi({\mathbf{x}}_{m})-\frac{1}{N}\sum_{n}\Phi({\mathbf{x}}_{n}^{\prime})\right\|^{2}$
in some potentially very high (even infinite) dimensional feature space $\Phi$ to estimate a squared MMD. Notably, the estimator depends on the two distributions only through examples and (unlike adversarial losses) does not require any optimization. This makes it worthwhile to consider as a candidate for our function $U$, especially in regimes where we have access to at least a small number of unrealistic examples. The basic idea is that we would fix a relatively large number of realistic examples and compare it to a small batch of examples we wish to test for realism. Support for this idea also comes from Amir and Weiss (2021) who have shown that MMD can be used to construct an effective full-reference perceptual metric [^1] A full-reference metric takes two images as arguments where a no-reference metric only has a single input. which agrees with human judgments in determining the similarity of pairs of images. To construct the metric, each image was treated as a distribution over small patches.
It remains unclear how to use MMD to quantify the realism of a single data point. For an image, one might compare features averaged over image patches to the features obtained from patches of a larger dataset, and similar ideas have showed some promise in image quality assessment (e.g., Mittal et al., 2013; Zhang et al., 2015). But the limitations of this approach are also clear as not all realistic images will have statistics representative of the entire data distribution.
A bigger concern perhaps is that the statistical power of MMD is known to drop quickly as the dimensionality of the problem increases (Ramdas et al., 2015), suggesting that we might need a very large number of examples if we want to identify defects in reasonably sized images or videos.
The MMD estimator above makes fewer assumptions than is necessary for us. In particular, it seems reasonable to assume access to $P$ (or a good approximation) both from a conceptual and a practical point of view, given the power of today’s generative models. By incorporating $P$ into our definition of realism, we can hope to quantify realism more efficiently. MMD leaves it to us to choose $\Phi$ and does not provide a clear mechanism for incorporating $P$.
@@ 4 Universal critics
A similar argument holds for continuous sample spaces (Li and Vitányi, 1997, Chapter 4.5). That is, there is a corresponding $S$ for continuous sample spaces which sums over measures, or lower semicomputable semimeasures [^3] A semimeasure is a measure which integrates to a value less than 1. $S$ itself is an example of a semimeasure, meaning $\sum_{\mathbf{x}}S({\mathbf{x}})<1$. This is due to the halting problem causing some $n$ to correspond to programs which never stop running. For these $n$, we set $Q_{n}({\mathbf{x}})=0$. to be precise. A measure is semicomputable if it can be approximated from below to arbitrary precision, that is, it is enough to be able to compute approximations of a measure for it to be included in the mixture $S$. For simplicity, we will focus on discrete spaces even though continuous spaces are relevant in practice if we want to optimize for realism.
@@
$-U({\mathbf{x}})$ $=\log\frac{\pi_{m}Q_{m}({\mathbf{x}})}{\sum_{n}\pi_{n}Q_{n}({\mathbf{x}})}-\log\pi_{m}$
@@ 4.1 Batched universal critics
since we must have $Q_{m}=Q$ for some $m$. For ease of notation, we also write $\pi_{Q}$ to refer to $\pi_{m}$. The inequality follows because the terms we dropped from the sum are all non-negative. What this sandwich bound implies is that our universal critic works well as a replacement for the optimal critic $T_{Q}$ (Eq. 7) if either the KL divergence between $Q$ and $P$ is large or the complexity of $Q$, $\log(1/\pi_{Q})$, is low. This agrees with our intuition for realism. In particular, we are more likely to accept an alternative explanation of the data if the explanation is simple, that is, if it can be described in a few words (or bits). A sequence of zeros (Examples 1 and 2) is easy to detect because it is cheap to describe (“always output 0”). While the critic $U$ depends on $P$, it is universal in the sense that it does not depend on $Q$.
@@
Example 4. Consider a distribution over natural images $P$ and a distribution $Q_{0}$ which assigns all its mass to a single flat image, $Q_{0}({\mathbf{x}}=0)=1$. Based on our bounds above, we should expect $U$ to detect $Q_{0}$ as unrealistic since its entropy is low (increasing its KL divergence) and it is cheap to describe, that is, $\log(1/\pi_{Q_{0}})$ is small for any reasonable coding scheme. (In contrast, using $-\log P({\mathbf{x}})$ instead of $U({\mathbf{x}})$ would fail to detect $Q_{0}$ since natural image distributions generally assign high probability to flat images.) Similarly, images of Gaussian white noise would be detected since their distribution is cheap to describe as independent copies of a simple distribution. Note from these examples that the complexity (or coding cost) $\log(1/\pi_{Q})$ of a distribution $Q$ is different from its entropy as low-complexity distributions can have both low or high entropy.
Example 5. As another example, consider a distribution which has memorized a training set of natural images, $Q_{\mathcal{D}}({\mathbf{x}})\propto\sum_{{\mathbf{x}}^{\prime}\in\mathcal{D}}\delta_{{\mathbf{x}}^{\prime}}({\mathbf{x}}).$ This distribution will remain undetected since its complexity is high. To describe $Q_{\mathcal{D}}$, we would have to encode every image in the training set $\mathcal{D}$. While this means that $U$ performs poorly as an approximation of KL divergence in this instance, it is in line with our intuitive notion of realism since we would also fail to tell a single example generated by $P$ from a single example of the training set (which we assume was generated by sampling from $P$). That is, we would consider training set images to be realistic and so would our universal critic.
We can improve the critic’s odds of detecting $Q$ by feeding it multiple examples. We define a batched universal critic as a critic of the form

### v4

@@ Abstract
perceptual quality, realism, neural compression, generative adversarial networks
@@ 1 Introduction
Potential applications of such functions are plentiful and include anomaly detection (Ruff et al., 2021), deepfake detection (Sha et al., 2023; Pondoc et al., 2023), generative model evaluation (Theis et al., 2016; Heusel et al., 2017; Borji, 2019), model distillation (van den Oord et al., 2018; Yin et al., 2023), neural compression (Ballé et al., 2021; Yang et al., 2023), computational photography (Fang et al., 2020), and computer graphics (Herzog et al., 2012; Reinhard et al., 2013; Poole et al., 2023). Unfortunately, their implementation is extremely challenging. Our ability to generate realistic data is rapidly improving (e.g., Dhariwal and Nichol, 2021) yet no reliable candidates or recipes for constructing $U$ exist in machine learning today. This is not for a lack of trying. While some progress has been made in the detection of unrealistic examples, the design of functions that are robust to optimization (for tasks involving generation) has been less successful. The latter problem is significantly harder because our function now not only has to detect a limited set of artefacts but has to anticipate any unrealistic examples an optimization might run into. Weaknesses in a function’s design often only make themselves known once we start optimizing (Ding et al., 2021). Complicating the matter is the fact that the optimization depends on $U$ itself.
To give a more concrete example of the kind of tasks we are interested in, consider the following loss which naturally comes up in lossy compression. If ${\mathbf{x}}=g({\mathbf{z}})$ is the output of a neural network, we may want to find a representation ${\mathbf{z}}$ such that
@@
In this paper we will take the view that ${\mathbf{x}}$ is realistic if it appears to have come about in a particular way, which is another way of saying that ${\mathbf{x}}$ is a plausible sample of a distribution $P$ capturing the data generating process. What is considered realistic therefore depends on $P$. If $P$ is a distribution over natural images then most photos would qualify as realistic. While an MNIST image (LeCun and Cortes, 2010) would not be considered a realistic example of a natural image, we would still consider it to be realistic if $P$ is the distibution of MNIST digits.
@@
What has been referred to as realism (e.g., Fan et al., 2018; Theis and Wagner, 2021; Careil et al., 2023) is also often referred to as perceptual quality (e.g., Blau and Michaeli, 2018; Fang et al., 2020; Salehkalaibar et al., 2023). It is therefore natural to wonder to what extent human perception should factor into its formalization. Our approach to defining realism is normative, that is, we consider how an idealized observer would judge realism. Similar to how Bayesian inference does not take inspiration from neuroscience but Bayesian decisions resemble human decisions (e.g., Knill and Pouget, 2004), we too can hope that human perception agrees with our definition of realism because it addresses a similar task as that faced by humans. In Section 4.4, we will further make the case that batched universal critics not only generalize no-reference metrics and divergences—which represent the prevalent ways of formalizing realism—but are also a better model of a human observer.
@@ 2 Probability and Typicality
In this section we review the two most common approaches to capture realism found in machine learning, namely those based on probability and typicality, and their failures. Similar failures of probability and typicality have been documented in the anomaly detection literature (e.g., Choi et al., 2019; Le Lan and Dinh, 2021; Osada et al., 2023) but are worth repeating as they continue to be a source of confusion.
@@ 2.1 Probability
Example 1 (Probability). Consider a computer program simulating a sequence of independent and nearly unbiased coin tosses, ${\mathbf{x}}^{N}=(x_{1},\dots,x_{N})$ with $P(x_{n}=1)=0.5+\varepsilon$ for some very small $\varepsilon>0$. For reasonably large $N$, we would expect the program to output a number of 1s which is close to N/2 and we would suspect a bug if the program outputs a sequence of only 1s, yet this is the most probable sequence.
Example 1 shows that maximizing $P({\mathbf{x}})$ can lead to unrealistic examples. It also shows that $P({\mathbf{x}})$ would not detect a bug which causes a program to only output 1s. If instead we count the number of 1s, $k=\sum_{n=1}^{N}x_{n}$, and measure the probability of $k$, this bug could be detected. Does this mean we only need to find the right set of features? By ignoring some aspects of the data, we risk not detecting unrealistic examples. We might therefore conclude that we simply need to test sufficiently many features. Unfortunately this approach also runs into trouble. Consider testing whether ${\mathbf{x}}$ has 1s in even places and 0s in odd places, ${\mathbf{x}}=\texttt{0101..01}$. The probability of this sequence is approximately $2^{-N}$ so that we would reject it with high confidence if we happen to observe it. However, since all sequences have roughly the same probability, we would reject every sequence as unrealistic if we tested all features identifying a specific sequence.
Using densities instead of probabilities introduces an additional challenge, namely that our answer now depends on the parametrization of the data. If $P$ is an exponential distribution with rate 1, say, then values of ${\mathbf{x}}$ close to zero seem preferable over larger values if judged by their density. But if we consider ${\mathbf{y}}=e^{-{\mathbf{x}}}$ instead, then all values of ${\mathbf{y}}$ would now be considered equally preferable.
@@ 2.2 Weak Typicality
and elements from this set are considered weakly typical. While other notions of typicality exist, weak typicality is the one most commonly encountered in the machine learning literature (Nalisnick et al., 2019b; Choi et al., 2019; Dieleman, 2020). The AEP implies that as $N$ increases, the probability that a randomly drawn sequence is contained in the typical set $A^{N}_{\delta}$ approaches 1 for any $\delta>0$. That is, a realistic sequence is likely to be typical. (While we have stated the AEP for sequences of independent and discrete random variables, generalizations to dependent and continuous sources exist and are well known; e.g., Algoet and Cover, 1988).
The above suggests that instead of expecting the probability to be large, we should expect realistic ${\mathbf{x}}$ to have negative log-probability close to the entropy—or the probability to be roughly $2^{-H[{\mathbf{x}}]}$, especially if ${\mathbf{x}}$ is high-dimensional. It therefore appears that $|-\log P({\mathbf{x}})-H[{\mathbf{x}}]|$ would be a good candidate for a measure of realism (Choi et al., 2019; Nalisnick et al., 2019b). Unfortunately, also this definition fails to quantify realism as the following examples demonstrate.
Example 2 (Typicality). Consider again a sequence of independent coin tosses. If the coin is unbiased, then the log-probability of any sequence is exactly the entropy, $-\log_{2}P({\mathbf{x}}^{N})=N$. In other words, in this case the probability of ${\mathbf{x}}^{N}$ under $P$ is completely uninformative and the typical set contains every sequence. Does this mean that every sequence of coin flips is realistic? Clearly, there is a sense in which the sequence 0000000000 is less realistic than 1100010100 which is not captured by weak typicality.
Example 3 (Typicality). As another example, consider a multivariate Gaussian distribution with density $p({\mathbf{x}})\propto\exp(-\|{\mathbf{x}}\|^{2})$. With high probability, the negative log-density of a random sample will be close to the differential entropy, which amounts to the norm $\|{\mathbf{x}}\|$ being roughly constant. While we would expect realistic examples from our distribution to look like uncorrelated noise, optimizing for typicality will only constrain the norm. If ${\mathbf{x}}$ represents an image, our optimization will merely adjust its contrast but will not decorrelate pixels as one might hope.
@@
An extended discussion of typicality can be found in Appendix A.
@@ 3 Divergences
More successful notions of realism are based on divergences between a ground-truth data distribution $P$ and a distribution $Q$ which we are trying to evaluate. In line with our intuitive notion of realism, if a divergence is zero, then instances of $Q$ are indistinguishable from instances of $P$, that is, we have perfect realism.
In coding theory, formalizing realism in terms of divergences (Matsumoto, 2018; Blau and Michaeli, 2019; Chen et al., 2022) has resulted in an improved understanding of the lossy compression problem and novel methods to solve them (e.g., Theis and Agustsson, 2021). In practical applications, generative adversarial networks (GANs; Goodfellow et al., 2014) trained with adversarial losses (which approximate divergences) significantly advanced the state of the art in the perceptual quality of generated images (e.g., Denton et al., 2015; Ledig et al., 2017). For the evaluation of generated images, the Fréchet inception distance (Heusel et al., 2017) has established itself as the method of choice and is based on a divergence between distributions over feature activations.
@@ 3.1 Adversarial Losses
Note that this optimal critic depends on the distribution $q$ that we are trying to evaluate. In contrast, in our setting we may only have access to a single instance or a few instances drawn from $q$. Furthermore, the dependence of the critic on $q$ is responsible for optimization instabilities that are known to plague adversarial training and which we would like to avoid. In Section 4 we will discuss critics which are universal in the sense that they do not depend on $q$ and therefore do not require adversarial training.
@@ 3.2 Maximum Mean Discrepancy
Maximum mean discrepancy (MMD; Gretton et al., 2012) refers to a class of divergences which have been used for hypothesis testing as well as for generating realistic images (Li et al., 2015; Dziugaite et al., 2015). Given two sets of i.i.d. examples—${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{M}$ and $\tilde{\mathbf{x}}_{1},\dots,\tilde{\mathbf{x}}_{N}$—estimates of MMD can be used to decide whether the two sets were drawn from the same distribution. Formally, we compute
$\text{MMD}^{2}({\mathbf{x}}^{M},\tilde{\mathbf{x}}^{N})=\textstyle\left\|\frac{1}{M}\sum_{m}\Phi({\mathbf{x}}_{m})-\frac{1}{N}\sum_{n}\Phi(\tilde{\mathbf{x}}_{n})\right\|^{2}$
in some potentially very high (even infinite) dimensional feature space $\Phi$ to estimate a squared MMD. Notably, the estimator depends on the two distributions only through examples and unlike adversarial losses does not require optimization of any critic. This makes it worthwhile to consider as a candidate for our function $U$, especially in regimes where we have access to at least a small number of unrealistic examples. The basic idea is that we would fix a relatively large number of realistic examples and compare it to a small batch of examples we wish to test for realism. Support for this idea also comes from Amir and Weiss (2021) who have shown that MMD can be used to construct an effective full-reference perceptual metric [^1] A full-reference metric takes two images as arguments where a no-reference metric only has a single input. which agrees with human judgments in determining the similarity of pairs of images. To construct the metric, each image was treated as a distribution over small patches.
It remains unclear how to use MMD to quantify the realism of a single data point without a reference. For an image, one might compare features averaged over image patches to the features obtained from patches of a larger dataset of images, and similar ideas have shown promise in image quality assessment (e.g., Mittal et al., 2013; Zhang et al., 2015). But the limitations of this approach are also clear as not all realistic images have statistics representative of the entire data distribution.
A bigger concern perhaps is that the statistical power of MMD can drop quickly as the dimensionality of the problem increases [^2] This assumes that the difficulty of the estimation problem remains constant, as measured by the KL divergence between the two distributions being tested. (Ramdas et al., 2015), suggesting that we might need a very large number of examples if we want to identify defects in reasonably sized images or videos.
The MMD estimator makes fewer assumptions than is necessary for us. In particular, it seems reasonable to assume access to $P$ (or a good approximation) both from a conceptual and a practical point of view, given the power of today’s generative models. By incorporating $P$ into our definition of realism, we can hope to quantify realism more efficiently. MMD leaves it to us to choose $\Phi$ and does not provide a clear mechanism for incorporating $P$.
@@ 4 Universal Critics
A similar argument holds for continuous sample spaces (Li and Vitányi, 1997, Chapter 4.5). That is, there is a corresponding $S$ for continuous sample spaces which sums over measures, or lower semicomputable semimeasures [^4] A semimeasure integrates to a value less or equal 1. $S$ itself is an example of a semimeasure with $\sum_{\mathbf{x}}S({\mathbf{x}})<1$. This is due to the halting problem causing some unknowable set of indices $n$ to correspond to programs which never stop running. For these $n$, we set $Q_{n}({\mathbf{x}})=0$. to be precise. A measure is semicomputable if it can be approximated from below to arbitrary precision, that is, it is enough to be able to compute approximations of a measure for it to be included in the mixture $S$. For simplicity, we will focus on discrete spaces even though continuous spaces are relevant in practice if we want to optimize for realism.
@@
$-U({\mathbf{x}})$ $=\log P({\mathbf{x}})-\log S({\mathbf{x}})$ (10)
$=\log\frac{\pi_{m}Q_{m}({\mathbf{x}})}{\sum_{n}\pi_{n}Q_{n}({\mathbf{x}})}-\log\pi_{m}$ (11)
$=\log\text{Pr}(m\mid{\mathbf{x}})-\log\pi_{m}$ (12)
@@ 4.1 Batched Universal Critics
since we must have $Q_{m}=Q$ for some $m$. For ease of notation, we also write $\pi_{Q}$ to refer to $\pi_{m}$. The inequality follows because the terms we dropped from the sum are all non-negative. What this sandwich bound implies is that our universal critic works well as a replacement for the optimal critic $T_{Q}$ (Eq. 7) if the complexity of $Q$, $\log(1/\pi_{Q})$, is low relative to the KL divergence between $Q$ and $P$. This agrees with our intuition for realism. In particular, we are more likely to accept an alternative explanation of the data if the explanation is simple, that is, if it can be described in a few words (or bits). A sequence of zeros (Examples 1 and 2) is easy to detect because it is cheap to describe (“always output 0”). While the critic $U$ depends on $P$, it is universal in the sense that it does not depend on $Q$.
Example 4 (Low Complexity). Consider a distribution over natural images $P$ and a distribution $Q_{0}$ which assigns all its mass to a single flat image, $Q_{0}({\mathbf{x}}=0)=1$. Based on our bounds above, we should expect $U$ to detect $Q_{0}$ as unrealistic since it is cheap to describe, that is, $\log(1/\pi_{Q_{0}})$ is small for any reasonable coding scheme. In contrast, using $-\log P({\mathbf{x}})$ instead of $U({\mathbf{x}})$ would fail to detect $Q_{0}$ since natural image distributions generally assign high probability to flat images. Similarly, images of Gaussian white noise would be detected since their distribution is cheap to describe as independent copies of a simple distribution.
Note from Example 4 that low-complexity distributions can have both low or high entropy, that is, the complexity (or coding cost) $\log(1/\pi_{Q})$ of a distribution $Q$ is different from its entropy.
Example 5 (High Complexity). As another example, consider a distribution which has memorized a training set of natural images, $Q_{\mathcal{D}}({\mathbf{x}})\propto\sum_{{\mathbf{x}}^{\prime}\in\mathcal{D}}\delta_{{\mathbf{x}}^{\prime}}({\mathbf{x}}).$ This distribution will remain undetected since its complexity is high. To describe $Q_{\mathcal{D}}$, we would have to encode every image in the training set $\mathcal{D}$. On the one hand, this means that $U$ may perform poorly as an approximation of the KL divergence between $Q_{\mathcal{D}}$ and $P$ (due to the loose lower bound, Eq. 17). On the other hand, this behavior is in line with our intuitive notion of realism since we would also fail to tell a single example generated by $P$ from a single example selected from the training set. Like the universal critic, we consider training set images to be realistic [^5] More concretely, we can say that an average training set image would be considered realistic in the sense that $\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{Q_{\mathcal{D}}}[U({\mathbf{x}})]]=\mathbb{E}_{P}[U({\mathbf{x}})]\leq 0$ (Eq. 14)..
@@
We can further improve the critic’s odds of detecting $Q$ by feeding it multiple independent examples. We define a batched universal critic as a critic of the form